if A,B ARE ANGLES IN THE FIRST QUADRANT,COSA =5/13 , SINB=4/5 FIND SIN(A-B)

Grade:9

1 Answers

Prasenjit Maitra
7 years ago
First find SINA from COSA & COSB fromSINB
SINA=√(1-COS²A)=√(1-(5/13)²)=12/13
COSB=√(1-SIN²B)=√(1-(4/5)²)=3/5
SIN(A-B)=SINACOSB-COSASINB=(12/13 * 3/5) – (5/13 * 4/5) = 16/65